# On-Device Self-Supervised Learning of Low-Latency Monocular Depth from Only Events

### v1

@@ Abstract
Event cameras provide low-latency perception for only milliwatts of power. This makes them highly suitable for resource-restricted, agile robots such as small flying drones. Self-supervised learning based on contrast maximization holds great potential for event-based robot vision, as it foregoes the need to high-frequency ground truth and allows for online learning in the robot’s operational environment. However, online, onboard learning raises the major challenge of achieving sufficient computational efficiency for real-time learning, while maintaining competitive visual perception performance. In this work, we improve the time and memory efficiency of the contrast maximization learning pipeline. Benchmarking experiments show that the proposed pipeline achieves competitive results with the state of the art on the task of depth estimation from events. Furthermore, we demonstrate the usability of the learned depth for obstacle avoidance through real-world flight experiments. Finally, we compare the performance of different combinations of pre-training and fine-tuning of the depth estimation networks, showing that on-board domain adaptation is feasible given a few minutes of flight.
@@ 1 Introduction
Event cameras capture per-pixel changes in brightness at microsecond resolution, while consuming only milliwatts of power [13]. This enables low-latency perception and decision making in environments even when undergoing large visual motion, on platforms that are constrained in power budget, such as small drones.
@@
To make full use of the temporal information in the event stream, the learning pipeline consisting of network architecture and loss function should also operate at high frequency [24]. Ground truth for variables like optical flow or depth is often only available at lower rates of 10-20 Hz [37, 14, 5]. While some datasets allow upsampling of ground truth to higher rates [37, 3], reaching the temporal resolution of an event camera might be difficult and come at the cost of high data rates. This holds back supervised learning at the short time scales perceivable with event cameras.
Figure 1: Self-supervised learning pipeline to estimate dense depth from event-camera data. We improve the time and memory efficiency of the pipeline for enabling online learning onboard of a small drone. This leads to better depth estimates and more successful obstacle avoidance behavior.
@@
Contrast maximization [11] allows self-supervised learning of optical flow, depth and ego-motion from events alone [38, 24, 27, 12]. Because no ground truth is needed, learning and prediction can be done at frequencies of 100 Hz [24] or even 200 Hz [25], with the only limit being the network’s ability to integrate temporal information long enough to capture sufficient motion for learning.
@@
A major advantage of self-supervised learning is that it foregoes the time- and pain-staking process of obtaining ground truth. This enables learning to scale up to enormous datasets. An additional, but typically less emphasized advantage is that self-supervised learning can in principle be performed in the operational environment of a robot or other edge device. Such online, self-supervised learning greatly reduces the need for generalization of the learned model, as training happens directly on data sampled from the test distribution [29]. For visual tasks such as monocular depth estimation, this is particularly important, as generalization of this perceptual capability to different environments than the training environment is notoriously difficult [31, 10]. Online, self-supervised learning does bring additional challenges. The first, main challenge is that not only the network but also the learning framework should be computationally efficient enough to run onboard.
@@
In this work, we improve the efficiency of the pipeline such that on-device learning of low-latency monocular depth and ego-motion becomes feasible. We demonstrate continual learning on-board a small flying drone, and show the usability of the resulting depth for obstacle avoidance. Furthermore, we investigate various combinations of pre-training and on-board learning. When trained on event camera datasets, our small recurrent network shows state-of-the-art depth estimation performance among self-supervised approaches. Our work scratches at the unused potential of on-board, online self-supervised learning, promising smaller reality gaps, leading to better performance.
@@ 2 Related work
2.1 (Self-/Un)supervised monocular depth estimation
Self-supervised and unsupervised learning of monocular depth has garnered significant attention since the early works that focused on joint depth-pose estimation for static scenes [36, 30, 16]. These foundational studies have spurred further advancements to handle more complex scenarios, such as dynamic scenes, by integrating optical flow estimation [26, 35], leveraging regularization techniques [21], or incorporating motion segmentation [28]. Additionally, several studies have explored learning camera parameters [17, 6], which is particularly relevant for on-device learning scenarios with unknown cameras.
Wang et al.[33] demonstrate that recurrent networks can enhance depth estimation by effectively utilizing information from multiple frames, resulting in more consistent depth scale predictions. Similarly, Bian et al. [2] introduce a loss term to encourage scale consistency, addressing a critical challenge in depth estimation. Achieving depth predictions with a consistent scale significantly enhances the stability of robot control systems relying on these depth estimates. By combining recurrent architectures and scale-consistent training, our approach aligns with these advancements, offering a robust solution for dynamic, on-device learning during real-world operation.
2.2 Contrast maximization to learn from events
@@
In our work, we adopt the approach outlined in [24], as it aligns with our goal of achieving high-frequency estimation using a neural network. This method is particularly well-suited for handling event trajectories that may exhibit nonlinear patterns, which are common in dynamic and complex scenarios. By using a neural network trained to extract optical flow from event data, we aim to ensure both flexibility and accuracy in environments where traditional linear assumptions might fail.
Moreover, this approach supports efficient inference, which is critical for real-time applications such as robotic navigation or on-device learning. The ability to accommodate nonlinear event trajectories further broadens its applicability, making it a robust choice for extracting motion information from event-based sensors in challenging conditions.
2.3 Self-supervised event-based depth estimation
Early works focused on jointly estimating depth and pose directly from events, employing either contrast maximization techniques [38] or photometric error methods based on event frames [34]. A notable contribution by Zhu et al. was their ability to estimate metric depth through the incorporation of a stereo loss term, enabling absolute depth recovery. These methods demonstrated the potential of event-based sensing for depth and pose estimation in static scenes.
More recent research has expanded these approaches to address dynamic scenes [15], where traditional static-scene assumptions do not hold, and developed more principled frameworks for model-based contrast maximization to jointly estimate depth, ego-motion, and optical flow [27]. These advancements represent a significant step toward understanding and processing complex, real-world motion using event data.
@@ 2.3 Self-supervised event-based depth estimation
2.4 On-device (learning of) depth estimation
Already in 2016, Lamers et al. [20] demonstrated the feasibility of learning depth estimation onboard a small flapping-wing drone. Although their approach did not produce dense depth maps, it proved effective for navigation, marking an early milestone in on-device learning for aerial robotics.
More recent advancements have focused on generating dense depth maps onboard drones. Liu et al.[22] developed a system to estimate depth from images for obstacle avoidance on a tiny quadrotor, leveraging recorded real-world datasets for training. Bhattacharya et al. [1], on the other hand, used a simulation-based approach to train depth estimation from events. They successfully transitioned to real-world applications by performing offline fine-tuning on real-world data, enabling effective obstacle avoidance in practice.
@@ 3.1 Optical flow from contrast maximization
Contrast maximization [11, 12] assumes events =$\mathcal{E}=\{e_{k}\}$ are triggered by motion, meaning that a warp $e_{k}=(t_{k},\bm{x}_{k},p_{k})\mapsto e^{\prime}_{k}=(t_{\text{ref}},\bm{x}^{\prime}_{k},p_{k})$ with the correct motion estimate will align it with other events triggered by the same portion of a moving edge, increasing the contrast of the image of warped events (IWE).
We follow the contrast maximization framework as in [24], which estimates optical flow for thin slices of the event stream using a recurrent architecture. By concatenating flows $\bm{u}_{i}$ as $\Delta\bm{x}_{k}=\sum_{i}(\Delta t_{i}\bm{u}_{i})(e_{k})$, events can be warped iteratively to neighbouring slices, with the correct flows leading to sharp IWEs at all reference times along the trajectory:
@@ 3.3 Optimizations for on-device learning
For efficient prediction, we make use of a relatively small, 430k-parameter network. It consists of a strided convolutional encoder, a ConvGRU recurrent bottleneck, and a two-tailed convolutional decoder for depth and ego-motion (more details in the supplementary material). While one network forward pass takes less than a millisecond on an NVIDIA RTX 4090, computing the contrast maximization loss and backpropagating the resulting gradients each take up more than 10 ms.
@@
To make on-device learning feasible, we have to improve the efficiency of the components that make up the loss computation and network update: i) warping all events in the accumulated set of events $\mathcal{E}$ using a sampled optical flow to all reference times $t_{\text{ref}}\in[0,T]$, ii) splatting them to the IWE at that $t_{\text{ref}}$, iii) computing the gradient with respect to the network parameters.
@@
Previous work [24] warped and splatted events in batches using PyTorch functions. This has multiple inefficiencies. Batching different amounts of events together leads to padding with zeros, resulting in wasteful computation and memory usage. This also goes for warping events that already went out of the image space (and therefore do not contribute to the loss anymore). Furthermore, PyTorch does not have an optimized implementation for bilinear splatting, meaning you have to implement it yourself. This results in extra computations and memory consumption for things like automatic differentiation.
Figure 2: Runtime and peak increase in memory consumption for different phases of computing the contrast maximization loss on an NVIDIA RTX 4090 and Jetson Orin NX (log-scale). Naive PyTorch processes all events in a for-loop. While batching events together improves a lot over this, parallel processing of all events in CUDA results in even bigger speedups with less memory consumed.
@@
The abovementioned issues can be resolved by considering that all events independently contribute to the loss because they are summed in the IWEs, and that we can therefore parallelize over all the accumulated events $\mathcal{E}$. We implement the functions to do so in CUDA, and connect them to PyTorch as an extension. As is shown in Fig. 2, the resulting improvements are, depending on the device, roughly 100x in terms of runtime. Common datasets like UZH-FPV [9] and MVSEC [37] have between 1k and 10k events per bin, well within the range of these speedups. Furthermore, looking at the peak delta memory, we can see that not needing padding allows much lower memory consumption when event data is highly sparse. Because we now provide the analytical gradients in the CUDA backward kernel, these do not have to be computed through automatic differentiation, leading to efficiency improvements for the backward pass.
@@ 3.4 Using depth for obstacle avoidance
3.5 Robot
The experimental setup as shown in Fig. 3 consists of a custom 5-inch quadrotor with a total weight of approximately 800 g, including all sensors, actuators, on-board compute, and battery. All algorithms are implemented to run entirely onboard, leveraging an NVIDIA Jetson Orin NX embedded GPU to obtain data from the event-camera, perform learning and estimate depth in real-time. Control commands (yaw rate) based on the predicted depth maps are sent to the flight controller, a Kakute H7 mini running the open-source autopilot software PX4.222https://px4.io/ Communication between PX4 and the Orin is done using ROS2 [23]. An MTF-01 optic flow sensor and rangefinder enable stable autonomous flight where only yaw rate is controlled by the depth pipeline.
To keep the event rate down (around 1 Mev/s), we only turn on every fourth pixel on the DVXplorer Micro, resulting in a 160x120 stream (instead of 640x480) for the same field-of-view. These events are accumulated into 20 ms windows and made into a frame for the network. The whole events-to-depth pipeline is running at approximately 30 Hz while learning, consuming on average around 15 W. We include a RealSense D435i depth camera for logging purposes only. To plot ground-truth flight trajectories, we record the drone’s position using a motion capture system.
@@ 4 Experiments
4.1 Benchmarks
Figure 4: Qualitative results of disparity predictions on the DSEC disparity benchmark. Images are for visualization only, as disparity estimation is event-based. The same color map is applied to the disparity values from the stereo- and supervised-learning-based method from Cho et al. [8] and ours for easy comparison.
We train our proposed network on the training sets (detailed training sequences provided in the supplementary materials) with a batch size of 8 and a constant learning rate of 1e-4 with the Adam optimizer for 50 epochs. The quantitative evaluations on MVSEC and DSEC are provided in Tab. 1 and Tab. 2.
@@ 4.1 Benchmarks
Table 2: Quantitative evaluation on DSEC disparity benchmark. Due to the lack of other SSL (self-supervised learning) methods on the leaderboard, we compare against two representative SL (supervised learning) methods.

### v2

@@ Abstract
Event cameras provide low-latency perception for only milliwatts of power. This makes them highly suitable for resource-restricted, agile robots such as small flying drones. Self-supervised learning based on contrast maximization holds great potential for event-based robot vision, as it foregoes the need for high-frequency ground truth and allows for online learning in the robot’s operational environment. However, online, on-board learning raises the major challenge of achieving sufficient computational efficiency for real-time learning, while maintaining competitive visual perception performance. In this work, we improve the time and memory efficiency of the contrast maximization pipeline, making on-device learning of low-latency monocular depth possible. We demonstrate that online learning on board a small drone yields more accurate depth estimates and more successful obstacle avoidance behavior compared to only pre-training. Benchmarking experiments show that the proposed pipeline is not only efficient, but also achieves state-of-the-art depth estimation performance among self-supervised approaches. Our work taps into the unused potential of online, on-device robot learning, promising smaller reality gaps and better performance.
Figure 1: Online, on-device learning allows robots to “train in their test environment”. We improve the time and memory efficiency of the self-supervised contrast maximization pipeline, such that on-board learning of monocular depth from event camera data becomes possible. When deployed on a small drone, online learning leads to better depth estimates and more successful obstacle avoidance behavior.
@@ 1 Introduction
Event cameras capture per-pixel brightness changes at microsecond resolution, while consuming only milliwatts of power [14]. This combination enables low-latency perception and decision-making on agile but resource-constrained platforms such as small drones.
@@
To make full use of the temporal information in the event stream, the learning pipeline consisting of network architecture and loss function should also operate at high frequency [28]. Ground truth optical flow or depth is often only available at lower rates of 10-20 Hz [44, 15, 5]. While some datasets allow upsampling of ground truth to higher rates [44, 3], reaching the temporal resolution of an event camera might be difficult and come at the cost of high data rates. This holds back supervised learning at the short time scales perceivable with event cameras.
@@
Contrast maximization [12] allows self-supervised learning (SSL) of optical flow, depth and ego-motion from events alone [45, 28, 32, 13]. Since no ground truth is needed, learning and prediction can run at higher frequencies of 100 Hz [28] or even 200 Hz [29], with the main limiting factor being the network’s ability to integrate information over increasingly sparse inputs as the rate increases.
@@
A major advantage of SSL is that it foregoes the costly process of obtaining ground truth, which enables learning to scale to large unlabeled datasets. An additional, but typically less emphasized advantage is that SSL can in principle be performed in the operational environment of a robot or other edge device. Such online SSL greatly reduces the need for generalization of the learned model, as training happens directly on data sampled from the test distribution [34]. For visual tasks like monocular depth estimation, this is particularly important, as generalization of this perceptual capability to environments different from the training environment is notoriously difficult [37, 11].
@@
Online SSL, however, introduces additional challenges. Chief among them is that not only the network but also the learning framework should be computationally efficient enough to run on board. In this work, we improve the efficiency of the contrast maximization pipeline such that on-device learning of low-latency monocular depth and ego-motion becomes feasible. We demonstrate continual learning of this complex visual task on board a small flying drone, and show the usability of the resulting depth for obstacle avoidance. Furthermore, we investigate various combinations of pre-training and on-board learning. When trained on event camera datasets, our small recurrent network shows state-of-the-art depth estimation performance among self-supervised approaches. Our work taps into the unused potential of on-board, online SSL, promising smaller reality gaps, leading to better performance.
@@ 2 Related work
SSL of monocular depth.
Self-supervised learning of monocular depth has garnered significant attention since the early works that focused on joint depth-pose estimation for static scenes [43, 35, 17]. These foundational studies have spurred further advancements to handle more complex scenarios, such as dynamic scenes, by integrating optical flow estimation [31, 42], leveraging regularization techniques [22], or incorporating motion segmentation [33]. Additionally, several works have explored learning camera parameters [18, 6], which is particularly relevant for on-device learning scenarios with unknown cameras.
Wang et al. [39] demonstrate that recurrent networks can enhance depth estimation by effectively utilizing information from multiple frames, resulting in more consistent depth scale predictions. Similarly, Bian et al. [2] introduce a loss term to encourage scale consistency, addressing a critical challenge in depth estimation. Achieving depth predictions with a consistent scale significantly enhances the stability of robot control systems relying on these depth estimates. By combining recurrent architectures and scale-consistent training, our approach aligns with these advancements, offering a robust solution for on-device learning during real-world operation.
SSL through contrast maximization.
@@
In our work, we adopt the approach outlined in [28], as it aligns with our goal of achieving high-frequency estimation using a neural network. Moreover, this approach supports efficient inference, which is critical for real-time applications such as robotic navigation or on-device learning. The ability to accommodate nonlinear event trajectories further broadens its applicability, making it a robust choice for extracting motion information in challenging conditions where traditional linear assumptions might fail.
SSL of depth from events.
Early works focused on jointly estimating depth and pose directly from events, employing either contrast maximization techniques [45] or photometric error methods based on event frames [41]. A notable contribution by Zhu et al. [45] was their ability to estimate metric depth through the incorporation of a stereo loss term, enabling absolute depth recovery. These methods demonstrated the potential of event-based sensing for depth and pose estimation in static scenes.
More recent research has expanded these approaches to address dynamic scenes [16], where traditional static-scene assumptions do not hold, and developed more principled frameworks for model-based contrast maximization to jointly estimate depth, ego-motion, and optical flow [32]. These advancements represent a significant step toward estimating complex, real-world motion using event data.
@@
Learning depth for drones.
Already in 2016, Lamers et al. [21] demonstrated the feasibility of learning depth estimation on board a small flapping-wing drone. Although their approach did not produce dense depth maps, it proved effective for navigation, marking an early milestone in on-device learning for aerial robotics. Several works [30, 23] combine unsupervised depth learning with an analytic odometry-flow pipeline (from external sources) to achieve metric monocular depth. While their networks are lightweight enough for embedded hardware, they are not actually used on drones.
Recent works have focused on generating dense depth maps on board drones. Liu et al. [24] developed a system to estimate depth from images for obstacle avoidance on a tiny quadrotor, leveraging recorded real-world datasets for training. Bhattacharya et al. [1], on the other hand, used a simulation-based approach to train depth estimation from events. They successfully transitioned to real-world applications by performing offline fine-tuning on real-world data, enabling effective obstacle avoidance in practice.
@@ 3.1 Optical flow from contrast maximization
Each pixel in an event camera independently detects changes in brightness and generates an event $e_{k}=(t_{k},\bm{x}_{k},p_{k})$ when this change exceeds a preset threshold. The polarity $p_{k}\in\{+1,-1\}$ indicates whether the brightness at pixel $\bm{x}_{k}$ and time $t_{k}$ increased or decreased. Contrast maximization [12, 13] assumes these events =$\mathcal{E}=\{e_{k}\}_{k=1}^{N_{e}}$ are triggered by motion, meaning that a warp $e_{k}=(t_{k},\bm{x}_{k},p_{k})\mapsto e^{\prime}_{k}=(t_{\text{ref}},\bm{x}^{\prime}_{k},p_{k})$ with the correct motion estimate $\Delta\bm{x}_{k}$ will align it with other events triggered by the same portion of a moving edge, increasing the contrast of the image of warped events (IWE).
We follow the contrast maximization framework as in [28], which estimates optical flow for thin slices of the event stream using a recurrent architecture. By concatenating flows $\bm{u}_{i}$ as $\Delta\bm{x}_{k}=\sum_{i}(\Delta t_{i}\bm{u}_{i})(e_{k})$, events can be warped iteratively to neighboring slices, with the correct flows leading to sharp IWEs at all reference times along the trajectory:
@@ 3.2 Combining depth and ego-motion into flow
Figure 2: Top left: events $e$ of different polarities are warped forward and backward by bilinearly sampled (BS) optical flows $\bm{u}$. Events warped outside the image are discarded. Next, events are bilinearly splatted to IWEs (images of warped events) at all reference times $t^{*}_{\text{ref}}$. Bottom left: batched processing of events such as in [28] requires zero-padding bins of events to equal length to facilitate simultaneous warping to neighboring reference times. In contrast, our per-event parallel processing in CUDA warps all events independently, doing away with padding and allowing to warp only those events still in the image space. Right: Runtime and peak increase in memory consumption for different phases of computing the contrast maximization loss on an NVIDIA RTX 4090 and Jetson Orin NX. 10% of each bin is made up of padding, which is not processed by the CUDA implementation. We indicate the range of events per bin for common datasets [9, 44] in black. Naive PyTorch processes all events in a for-loop. While batching events together improves a lot over this, parallel processing of all events in CUDA results in even bigger speedups with less memory consumed.
@@ 3.3 Optimizations for on-device learning
For efficient prediction, we make use of a lightweight, 430k-parameter network inspired by [40]. It consists of a strided convolutional encoder, a ConvGRU recurrent bottleneck, and a two-tailed convolutional decoder for depth and ego-motion (more details in the supplementary material). While one network forward pass takes less than a millisecond on an NVIDIA RTX 4090, computing the contrast maximization loss and backpropagating the resulting gradients each take up more than 10 ms.
@@
To make on-device learning feasible, we have to improve the efficiency of the components that make up the loss computation and network update: i) warping all events in the accumulated set of events $\mathcal{E}$ using a sampled optical flow to all reference times $t_{\text{ref}}\in[0,T]$, ii) bilinearly splatting them to the IWE at that $t_{\text{ref}}$, iii) computing the gradient with respect to the network parameters.
@@
Previous work [28] warped and splatted events in batches using PyTorch functions. This has multiple inefficiencies. Batching different amounts of events together leads to padding with zeros, resulting in wasteful computation and memory usage. This also goes for warping events that already went out of the image space (and therefore do not contribute to the loss anymore). Furthermore, some operations (like bilinear splatting) do not have optimized implementations in PyTorch. All this combined results in extra computational and memory overhead due to intermediate tensor allocations, multiple instead of single kernels, computation graph tracking and scattered memory accesses.
The abovementioned issues can be resolved by considering that all events independently contribute to the loss since they are summed in the IWEs, and that we can therefore parallelize over all the accumulated events $\mathcal{E}$. We implement the functions to do so in CUDA, getting rid of most of the overhead, and connect them to PyTorch as an extension.
Specifically, we assign one CUDA thread to handle one or more events in parallel, read off their positions in the 3D domain $(x,y,t)$, and then apply the flow fields to “push” each event through time either forward or backward in an iterative fashion. During the backward pass, we use the stored warped-point positions along each time step (forward and backward) to compute partial derivatives w.r.t the flow vector $\frac{\partial\mathcal{L}}{\partial\mathbf{f}}$. As in the forward pass, we rely on bilinear interpolation weights at each $(x,y)$ to distribute gradients to the four nearest flow-vector cells. By avoiding padding and unnecessary warping computations, we achieve significantly better parallelization on GPUs.
@@
As shown in Fig. 2, the resulting improvements are, depending on the device, roughly 100x in terms of runtime, and 2-5x in terms of memory consumption. Common datasets like UZH-FPV [9] and MVSEC [44] have between 1k and 10k events per bin, well within the range of these speedups. Furthermore, looking at the peak delta memory, the removal of padding yields much lower memory consumption when event data is highly sparse. Because we now provide the analytical gradients in the CUDA backward kernel, these do not have to be computed through automatic differentiation, leading to further efficiency improvements.
@@ 3.4 Using depth for obstacle avoidance
3.5 Drone and flight environment
The experimental setup as shown in Fig. 3 consists of a custom 5-inch quadrotor with a total weight of approximately 800 g, including all sensors, actuators, on-board compute and battery. All algorithms are implemented to run entirely on board, using an NVIDIA Jetson Orin NX embedded GPU to receive data from the event camera, perform learning and estimate depth in real time. Control commands (yaw rate) based on the predicted depth maps are sent to the flight controller, a Kakute H7 Mini running the open-source autopilot software PX4. Communication between PX4 and the Orin is done using ROS2 [25]. An MTF-01 optic flow sensor and rangefinder enables stable autonomous flight where only yaw rate is controlled based on the depth estimate.
To keep the event rate down (below 1 Mev/s), we only turn on every fourth pixel on the DVXplorer Micro, resulting in a 160x120 stream (instead of 640x480) for the same field-of-view. These events are accumulated into 20 ms windows and made into a frame for the network. The whole events-to-depth pipeline is running at approximately 30 Hz while learning, consuming on average around 9 W. We include a RealSense D435i depth camera for logging purposes only. To plot ground-truth flight trajectories, we record the drone’s position using a motion capture system.
@@ 4 Experiments
4.1 Event-based depth benchmarks
Figure 4: Qualitative results of disparity predictions on the DSEC disparity benchmark. Images are for visualization only, as disparity estimation is event-based. The same color map is applied to the disparity values from the stereo- and supervised-learning-based method from Cho et al. [8] and our monocular, self-supervised learning method for easy comparison.
Setup.
We train our proposed network on the training sets with a batch size of 8 and a constant learning rate of 1e-4 with the Adam optimizer for 50 epochs (more details in the supplementary material). We do truncated backpropagation through time, with a backward pass/gradient update conducted every 10 forward passes. Detaching the network while not resetting its state ensures bounded memory usage, mitigates potential gradient explosion/vanishing, and allows the network to retain temporal context effectively. The quantitative evaluations on MVSEC and DSEC are provided in Tab. 1 and Tab. 2.
@@ 4.1 Event-based depth benchmarks
Table 2: Quantitative evaluation on the DSEC disparity benchmark. Due to the lack of other monocular SSL (self-supervised learning) methods on the leaderboard, we compare against two representative stereo-based SL (supervised learning) methods.
Results.